OpenStudy (lgbasallote):

2+7+8 = 141623 9+4+3 = 362759 5+4+5 = 202541 8+3+7 = 245677 6+3+9 = ? easy one >:))

OpenStudy (lgbasallote):

@FoolForMath

hero (hero):

2*8 = 16 2*7 = 14 16 + 14 - 7 = 23

OpenStudy (lgbasallote):

haha nice one =)))

OpenStudy (anonymous):

Genius.

OpenStudy (anonymous):

185469.

OpenStudy (anonymous):

Nice, hero.

OpenStudy (lgbasallote):

didnt know hero was as good as ffm =))))

hero (hero):

Like you said, it was an easy one.

OpenStudy (anonymous):

Okay I am the hungry one. See you guys later :)